# From Policy to Logic for Efficient and Interpretable Coverage Assessment

### v1

@@ Introduction
Recent advances in Large Language Models (LLMs) have demonstrated considerable promise in decision-making across domains such as legal analysis and healthcare policy interpretation (ryu2025divide; pan-etal-2023-logic; xu-etal-2024-faithful), owing to their ability to process and interpret complex natural language. While it is generally accepted that LLMs are adept at analyzing textual information, they often face difficulties when analyzing lengthy complex texts. Effective reasoning in these contexts requires models to explicitly reference factual details and policy language, which are frequently dispersed throughout the document. Furthermore, LLMs may exhibit hallucinations and inconsistencies in their reasoning (LegalfictionsDahletal2024).
@@
Chain-of-thought (CoT) prompting (wei2022chain) is a commonly explored technique (kant2025towards) for guiding LLMs through multi-step reasoning processes. However, CoT approaches are not immune to lack of interpretability, generating inconsistent reasoning, and can be computationally expensive when applied at scale.
Expert systems have been used to simulate the reasoning abilities of human experts by encoding domain-specific knowledge and reasoning processes (li2024mediq; garrido2025gofai). These systems operate on structured representations of knowledge, such as facts and “if-then” rules, using deterministic logic to systematically process knowledge. By applying predefined rules, it enables expert systems ensure consistency and interpretability.
@@
The use of LLMs to translate natural language into formal rules has gained significant attention in recent research. Techniques such as prompting with CoT and neuro-symbolic approaches (wei2022chain; nezhad2025symcode; pmlr-v284-nezhad25a) have shown promising results in this area. Contractual language often encodes logical relationships in natural language that humans interpret with ease. Several recent systems, including CLOVER (ryu2025divide), LOGIC-LM (pan-etal-2023-logic), LogicLLaMA (yang-etal-2024-harnessing), ProSLM (vakharia2024proslm), Thought Like Pro (tan2024thought), SymbCoT (xu-etal-2024-faithful), and LLM-Tres (toroghi-etal-2024-verifiable), have explored methods for translating natural language into first-order logic representations. These approaches aim to bridge the gap between unstructured contract language and machine-interpretable rules. Our work is in line with these efforts to formalize natural language reasoning, but differs in that we employ a rule-based expert system to operationalize domain knowledge through executable reasoning rather than formal logical translation.
@@
LLMs have demonstrated notable progress in supporting legal decision-making tasks, including the following examples. guha2023legalbench introduced LegalBench, a comprehensive suite of benchmarks to evaluate the legal reasoning capabilities of LLMs. yao-etal-2025-elevating proposed a reinforcement learning-based approach for legal question answering. mishra-etal-2025-investigating conducted an error analysis of LLM reasoning in civil procedure contexts. LegalfictionsDahletal2024 examined the prevalence of hallucinations in LLMs applied to legal domains. Shengbin2024lawllm presented LawLLM, a system offering legal reasoning services through fine-tuning techniques. shen2025law introduced a reasoning schema for legal tasks that integrates factual grounding. Similarly, shi2025legalreasoner proposed LegalReasoner, which first identifies disputes to decompose complex cases and then performs step-wise reasoning. Our approach focuses on using actionable rules generated from text.
@@
cummins2025computable explored the potential of using computable rules for decision-making in a specific domain, leveraging Prolog-like logical representations to encode policy logic in a machine-readable format. Their work highlights the importance of formal rule structures in enabling automated reasoning, but relies heavily on manual encoding and domain expertise for rule construction.
Building on this foundation, kant2025towards conducted an extensive evaluation of the current generation of LLMs in producing structured rules from policy documents. In their study, the authors systematically compared the decision-making capabilities of LLMs using both prompt-based approaches and rule-based frameworks. LLMs were tasked with generating structured rules from policy text, guided by human-designed schemas featuring specific attributes and helper functions. Their findings show that most LLMs exhibit markedly improved reasoning performance when integrated with explicit rule-based scaffolding, underscoring the value of structured logic in complex adjudication tasks. However, their approach is limited by the scope of the provided schemas and helper functions, constraining reasoning to facts explicitly represented in the input.
@@
This formulation, functionally equivalent to contrastive objectives like InfoNCE (oord2018representation), forces the model to distinguish the single governing passage from a set of highly relevant but non-dispositive distractors from the same document. Queries are constructed as:
@@ Results
Our error analysis across all procedure codes and seven plans reveals that most rule failures can be attributed to two main factors. The first and most prevalent is that the correct attribute is sometimes not incorporated into the rule creation process. This accounts for 73.5% of incorrect cases, where a rule is triggered but does not represent the correct logic. This often occurs when the attribute list is lengthy, leading the model to overlook or “forget” attributes that appear later in the input sequence, a phenomenon also observed by (agrawal-etal-2024-evaluating; tao2025lost). The second failure mode, making up the remaining 26.5% of errors, involves instances where the model fails to generate a sufficient set of rules for certain samples, resulting in incomplete or missing logic and, thus, no rule being triggered for accurate coverage decisions. These findings underscore the importance of careful attribute selection and prompt engineering when constructing symbolic rules from complex policy language. Improving attribute coverage and accuracy is the most impactful path forward.

### v2

@@ Introduction
Recent advances in Large Language Models (LLMs) have demonstrated considerable promise in decision-making across domains such as legal analysis and healthcare policy interpretation (Ryu et al. 2025; Pan et al. 2023; Xu et al. 2024), owing to their ability to process and interpret complex natural language. While it is generally accepted that LLMs are adept at analyzing textual information, they often face difficulties when analyzing lengthy complex texts. Effective reasoning in these contexts requires models to explicitly reference factual details and policy language, which are frequently dispersed throughout the document. Furthermore, LLMs may exhibit hallucinations and inconsistencies in their reasoning (Dahl et al. 2024).
@@
Chain-of-thought (CoT) prompting (Wei et al. 2022) is a commonly explored technique (Kant et al. 2025) for guiding LLMs through multi-step reasoning processes. However, CoT approaches are not immune to lack of interpretability, generating inconsistent reasoning, and can be computationally expensive when applied at scale.
Expert systems have been used to simulate the reasoning abilities of human experts by encoding domain-specific knowledge and reasoning processes (Li et al. 2024; Garrido-Merchán and Puente 2025). These systems operate on structured representations of knowledge, such as facts and “if-then” rules, using deterministic logic to systematically process knowledge. By applying predefined rules, it enables expert systems ensure consistency and interpretability.
@@
The use of LLMs to translate natural language into formal rules has gained significant attention in recent research. Techniques such as prompting with CoT and neuro-symbolic approaches (Wei et al. 2022; Nezhad et al. 2025; Nezhad and Agrawal 2025) have shown promising results in this area. Contractual language often encodes logical relationships in natural language that humans interpret with ease. Several recent systems, including CLOVER (Ryu et al. 2025), LOGIC-LM (Pan et al. 2023), LogicLLaMA (Yang et al. 2024), ProSLM (Vakharia et al. 2024), Thought Like Pro (Tan et al. 2024), SymbCoT (Xu et al. 2024), and LLM-Tres (Toroghi et al. 2024), have explored methods for translating natural language into first-order logic representations. These approaches aim to bridge the gap between unstructured contract language and machine-interpretable rules. Our work is in line with these efforts to formalize natural language reasoning, but differs in that we employ a rule-based expert system to operationalize domain knowledge through executable reasoning rather than formal logical translation.
@@
LLMs have demonstrated notable progress in supporting legal decision-making tasks, including the following examples. Guha et al. (2023) introduced LegalBench, a comprehensive suite of benchmarks to evaluate the legal reasoning capabilities of LLMs. Yao et al. (2025) proposed a reinforcement learning-based approach for legal question answering. Mishra et al. (2025) conducted an error analysis of LLM reasoning in civil procedure contexts. Dahl et al. (2024) examined the prevalence of hallucinations in LLMs applied to legal domains. Yue et al. (2024) presented LawLLM, a system offering legal reasoning services through fine-tuning techniques. Shen et al. (2025) introduced a reasoning schema for legal tasks that integrates factual grounding. Similarly, Shi et al. (2025) proposed LegalReasoner, which first identifies disputes to decompose complex cases and then performs step-wise reasoning. Our approach focuses on using actionable rules generated from text.
@@
Cummins et al. (2025) explored the potential of using computable rules for decision-making in a specific domain, leveraging Prolog-like logical representations to encode policy logic in a machine-readable format. Their work highlights the importance of formal rule structures in enabling automated reasoning, but relies heavily on manual encoding and domain expertise for rule construction.
Building on this foundation, Kant et al. (2025) conducted an extensive evaluation of the current generation of LLMs in producing structured rules from policy documents. In their study, the authors systematically compared the decision-making capabilities of LLMs using both prompt-based approaches and rule-based frameworks. LLMs were tasked with generating structured rules from policy text, guided by human-designed schemas featuring specific attributes and helper functions. Their findings show that most LLMs exhibit markedly improved reasoning performance when integrated with explicit rule-based scaffolding, underscoring the value of structured logic in complex adjudication tasks. However, their approach is limited by the scope of the provided schemas and helper functions, constraining reasoning to facts explicitly represented in the input.
@@
This formulation, functionally equivalent to contrastive objectives like InfoNCE (Oord et al. 2018), forces the model to distinguish the single governing passage from a set of highly relevant but non-dispositive distractors from the same document. Queries are constructed as:
@@ Results
Our error analysis across all procedure codes and seven plans reveals that most rule failures can be attributed to two main factors. The first and most prevalent is that the correct attribute is sometimes not incorporated into the rule creation process. This accounts for 73.5% of incorrect cases, where a rule is triggered but does not represent the correct logic. This often occurs when the attribute list is lengthy, leading the model to overlook or “forget” attributes that appear later in the input sequence, a phenomenon also observed by (Agrawal et al. 2024; Tao et al. 2025). The second failure mode, making up the remaining 26.5% of errors, involves instances where the model fails to generate a sufficient set of rules for certain samples, resulting in incomplete or missing logic and, thus, no rule being triggered for accurate coverage decisions. These findings underscore the importance of careful attribute selection and prompt engineering when constructing symbolic rules from complex policy language. Improving attribute coverage and accuracy is the most impactful path forward.